# Street Gaussians for Modeling Dynamic Urban Scenes

### v1

@@ Abstract
This paper aims to tackle the problem of modeling dynamic urban street scenes from monocular videos. Recent methods extend NeRF by incorporating tracked vehicle poses to animate vehicles, enabling photo-realistic view synthesis of dynamic urban street scenes. However, significant limitations are their slow training and rendering speed, coupled with the critical need for high precision in tracked vehicle poses. We introduce Street Gaussians, a new explicit scene representation that tackles all these limitations. Specifically, the dynamic urban street is represented as a set of point clouds equipped with semantic logits and 3D Gaussians, each associated with either a foreground vehicle or the background. To model the dynamics of foreground object vehicles, each object point cloud is optimized with optimizable tracked poses, along with a dynamic spherical harmonics model for the dynamic appearance. The explicit representation allows easy composition of object vehicles and background, which in turn allows for scene editing operations and rendering at 133 FPS (1066$\times$1600 resolution) within half an hour of training. The proposed method is evaluated on multiple challenging benchmarks, including KITTI and Waymo Open datasets. Experiments show that the proposed method consistently outperforms state-of-the-art methods across all datasets. Furthermore, the proposed representation delivers performance on par with that achieved using precise ground-truth poses, despite relying only on poses from an off-the-shelf tracker. The code is available at https://zju3dv.github.io/street_gaussians/.
@@ 1 Introduction
Figure 1: Rendering results on the Waymo dataset [44]. With only estimated trajectories, our method produces high-quality rendering at 133 FPS (1066$\times$1600) within half an hour of training. Current SOTA method [55] requires accurate ground-truth trajectories, and has low training and rendering speed.
@@
With the development of neural scene representations, there have been some methods [27, 45, 37, 34, 69] that attempt to reconstruct street scenes with neural radiance fields [30]. To improve the modeling capability, Block-NeRF [45] divides the target street into several blocks and represents each one with a NeRF network. Although this strategy enables photo-realistic rendering of large-scale street scenes, Block-NeRF suffers from long training time due to the large amount of network parameters. Moreover, it cannot handle dynamic vehicles on the street, which are crucial aspects in autonomous driving environment simulation.
@@
Recently, some methods [33, 62, 55, 19] propose to represent dynamic street scenes as compositional neural representations that consist of foreground moving cars and static background. To handle the dynamic car, they leverage tracked vehicle poses to establish the mapping between the observation space and the canonical space, where they use NeRF networks to model the car’s geometry and appearance. As a result, these methods are sensitive to the accuracy of tracked bounding boxes. In addition, they are still limited to the high training cost and low rendering speed.
@@
In this work, we propose a novel explicit scene representation for reconstructing dynamic 3D street scenes from images. The basic idea is utilizing point clouds to build dynamic scenes, which significantly increases the training and rendering efficiency while decreasing the dependency on the accuracy of tracked vehicle poses. Specifically, we decompose urban street scenes into the static background and moving vehicles, which are separately built based on 3D Gaussians [16]. To handle the dynamics of foreground vehicles, we model their geometry as a set of points with optimizable tracked vehicle poses, where each point stores learnable 3D Gaussian parameters. Furthermore, the time-varying appearance is represented by a 4D spherical harmonics model that uses a time series function to predict spherical harmonics coefficients at any time step. Thanks to the dynamic Gaussian representation, we can faithfully reconstruct the target urban street within half an hour and achieve real-time rendering (133FPS@1066x1600).
Building upon the proposed scene representation, we further develop a tracked pose optimization strategy. The optimizable input tracked poses ensure a better alignment between the rendered and input videos. Experiments reveal that our method yields comparable results to that achieved with ground-truth poses, while only utilizing poses from an off-the-shelf tracker [53], thanks to the better gradient propagation facilitated by our explicit representation.
@@
We evaluate the proposed method on Waymo Open [44] (Waymo) and KITTI [13] datasets, which present dynamic street scenes with complex vehicle motions and various environment conditions. Across all datasets, our approach achieves state-of-the-art performance in terms of rendering quality, while being rendered over 100 times faster than previous methods [55, 33]. Furthermore, detailed ablations and scene editing applications are conducted to demonstrate the effectiveness of proposed components and the flexibility of the proposed representation, respectively.
@@
  • We develop a tracked pose optimization strategy along with a 4D spherical harmonics appearance model to handle the dynamics of moving vehicles.
@@
2 Related work
Simulation environments for autonomous driving.  Existing self-driving simulation engines such as CARLA [9] or AirSim [40] suffer from costly manual effort to create virtual environments and the lack of realism in the generated data. In recent years, a lot of effort has been put into building sensor simulations from autonomous driving data captured in real scenes. Some works [29, 61, 10] concentrate on LiDAR simulation by aggregating LiDAR and reconstructing textured primitives. However, they have difficulty handling high-resolution images and usually produce noisy appearance. Other works [64, 7, 49] reconstruct objects from multi-view images and LiDAR input, which can be interacted with other environments. However, these methods are restricted to existing images and fail to render novel views. Some methods utilize neural fields to perform multiply tasks including view synthesis [62, 33, 15], perception [19, 12, 69], generation [41, 32, 57, 60, 21] and inverse rendering [50] on driving scenes. However, they struggle with long training time and slow rendering speed. In contrast, Our method focuses on performing real-time rendering of dynamic urban street scenes, which is crucial for autonomous driving simulation.
Static scene modeling.  Neural scene representation proposes to represent 3D scenes with neural networks, which can model complex scenes from images through differentiable rendering. NeRF [30, 3, 4, 5, 31] represents continuous volumetric scenes with MLP networks and achieves impressive rendering results. Some works have been proposed to extend NeRF to urban scenes [37, 45, 14, 27, 26, 34, 47]. DNMP [27] models the scene with deformable mesh primitives initialized by voxelizing point clouds. NeuRas [26] takes scaffold mesh as input and optimizes the neural texture field to perform fast rasterization. Point-based rendering works [38, 18, 1, 8, 23] define learned neural descriptors on point clouds and perform differentiable rasterization with a neural renderer. However, they require dense point clouds as input and generate blurry results under regions with low point counts. A very recent work 3D Gaussian Splatting (3D GS) [16] defines a set of anisotropic Gaussians in 3D world and performs adaptive density control to achieve high-quality rendering results with only sparse point clouds input. However, 3D GS assumes the scene to be static and can not model dynamic moving objects.
Dynamic scene modeling.  Recent methods build 4D neural scene representation on single-object scenes [42, 35, 11, 22, 2, 24, 36, 25]. Some works learn a scene decomposition of outdoor scenes under the supervision of optical flow [48] or vision transformer feature [59]. However, their scene decomposition can not be edited, limiting the applications for autonomous driving simulation. Another line of works model the scene as the composition of moving object models and a background model [33, 19, 55, 62, 56] with neural fields, which is most similar to us. However, they require accurate object trajectories and suffer from high memory cost and slow inference speed. Point-based dynamic scene rendering is also investigated recently [58, 67]. Concurrent to our work [63, 65, 52, 28] also extend 3D GS to dynamic scenes, but they build their models on small-scale data while we focus on large-scale urban street scenes.
3 Method
Given a sequence of images captured from a moving vehicle in an urban street scene, our goal is to develop a model capable of generating photorealistic images for any given input time step and any viewpoints. Towards this objective, we propose a novel scene representation, named Street Gaussians, specifically designed for representing dynamic street scenes. As shown in the Figure 2, we represent a dynamic urban street scene as a set of point clouds, each corresponding to either the static background or a moving vehicle (Section 3.1). The explicit point-based representation allows easy composition of separate models, enabling real-time rendering as well as the decomposition of foreground objects for editing applications (Section 3.2). The proposed scene representation can be effectively trained using only RGB images, along with tracked vehicle poses from an off-the-shelf tracker, enhanced by our tracked vehicle pose optimization strategy (Section 3.3).
@@ 3.1 Street Gaussians
Background model.
The background model is represented as a set of points in the world coordinate system. Each point is assigned with a 3D Gaussian to softly represent the continuous scene geometry and color. The Gaussian parameters consist of a covariance matrix $\bm{\Sigma}_{b}$ and a position vector $\bm{\mu}_{b}\in\mathbb{R}^{3}$, which denotes the mean value. To avoid invalid covariance matrix during optimization, each covariance matrix is further reduced to a scaling matrix $\mathbf{S}_{b}$ and a rotation matrix $\mathbf{R}_{b}$, where $\mathbf{S}_{b}$ is characterized by its diagonal elements. and $\mathbf{R}_{b}$ is converted into a unit quaternion. The covariance matrix $\bm{\Sigma}_{b}$ can be recovered from $\mathbf{S}_{b}$ and $\mathbf{R}_{b}$ as:
@@
Object model.
Consider a scene containing $N$ moving foreground object vehicles. Each object is represented with a set of optimizable tracked vehicle poses and a point cloud, where each point is assigned a 3D Gaussian, semantic logits, and a dynamic appearance model.
@@
$\mathbf{R}_{w}$ $=\mathbf{R}_{o}\mathbf{R}_{t}^{T},$
@@
The semantic representation of the object model is different from that of the background model. The main difference is that the semantic of the object model is a one-dimensional scalar $\beta_{o}$ instead of a $M$-dimensional vector $\bm{\beta}_{b}$ like the background model. This semantic model for the foreground object vehicle model can be considered as a binary classification or a confidence prediction problem because there are only two semantic categories for the object, i.e., vehicle semantic class (from the tracker) and non-vehicle.
@@ 3.2 Rendering of Street Gaussians
To render Street Gaussians, we need to aggregate the contribution of each model to render the final image. Previous methods [33, 19, 62, 55] require compositional rendering with complex raymarching because of neural field representation. Instead, Street Gaussians can be rendered by contacting all the point clouds and projecting them to 2D image space. Specifically, given a rendered time step $t$, we first compute spherical harmonics Eq. 3 , and transform the object point cloud into the world coordinate system using and Eq. 2 according to tracked vehicle pose $(\mathbf{R}_{t},\mathbf{T}_{t})$. Then we concatenate the background point cloud and the transformed object point clouds to form a new point cloud. To project this point cloud to 2D image space with camera extrinsic $\mathbf{W}$ and intrinsic $\mathbf{K}$, we compute the 2D Gaussian for each point in the point cloud:
@@
where $\mathbf{J}$ is the Jacobian matrix of $\mathbf{K}$. $\mu^{\prime}$ and $\Sigma^{\prime}$ are the position and covariance matrix in 2D image space, respectively. Point-based $\alpha$-blending for each pixel is used to compute the color $\mathbf{c}$:
@@
Here $\alpha_{i}$ is the opacity $\alpha$ multiplied by the probability of the 2D Gaussian [66] and $\mathbf{c}_{i}$ is the color computed from spherical harmonics $\mathbf{z}$ with the view direction.
The semantic map can be rendered simply by changing color $c$ in Eq 5 to semantic logits $\bm{\beta}$.
@@ 3.3 Training
Tracking pose optimization.
Positions and covariance matrices of the object Gaussians during rendering in Section 3.2 are closely correlated with the tracked pose parameters as shown in Eq 2. However, bounding boxes produced by the tracker model are generally noisy. Directly using them to optimize our scene representation leads to degradation in rendering quality. As a result, we treat tracked poses as learnable parameters by adding a learnable transformation to each transformation matrix. Specifically, $\mathbf{R}_{t}$ and $\mathbf{T}_{t}$ in Eq 2 are replaced by $\mathbf{R}_{t}^{{}^{\prime}}$ and $\mathbf{T}_{t}^{{}^{\prime}}$ which are defined as:
@@
Loss function.
We jointly optimize our scene representation and tracked poses using the following loss function:
$\mathcal{L}=\mathcal{L}_{\text{color}}+\lambda_{1}\mathcal{L}_{\text{sem}}+\lambda_{2}\mathcal{L}_{\text{reg}}.$ (7)
In Eq 7, $\mathcal{L}_{\text{color}}$ is the reconstruction loss between rendered and observed images. $\mathcal{L}_{\text{sem}}$ is an optional per-pixel softmax-cross-entropy loss between rendered semantic logits and input 2D semantic segmentation predictions [20]. In order to prevent noisy input semantic labels from influencing the geometry, we stop the gradients of $\mathcal{L}_{\text{sem}}$ back to $\alpha$ in Eq 5. $\mathcal{L}_{\text{reg}}$ is an entropy regularization term used to remove floaters and enhance decomposition effects. Please refer to the supplementary material for the details. In practice, we set $\lambda_{1}=0.1$ and $\lambda_{2}=0.1$.

### v2

# Street Gaussians: Modeling Dynamic Urban Scenes with Gaussian Splatting

@@ Abstract
This paper aims to tackle the problem of modeling dynamic urban streets for autonomous driving scenes. Recent methods extend NeRF by incorporating tracked vehicle poses to animate vehicles, enabling photo-realistic view synthesis of dynamic urban street scenes. However, significant limitations are their slow training and rendering speed. We introduce Street Gaussians, a new explicit scene representation that tackles these limitations. Specifically, the dynamic urban scene is represented as a set of point clouds equipped with semantic logits and 3D Gaussians, each associated with either a foreground vehicle or the background. To model the dynamics of foreground object vehicles, each object point cloud is optimized with optimizable tracked poses, along with a 4D spherical harmonics model for the dynamic appearance. The explicit representation allows easy composition of object vehicles and background, which in turn allows for scene editing operations and rendering at 135 FPS (1066 * 1600 resolution) within half an hour of training. The proposed method is evaluated on multiple challenging benchmarks, including KITTI and Waymo Open datasets. Experiments show that the proposed method consistently outperforms state-of-the-art methods across all datasets. The code will be released to ensure reproducibility. $\times$
Keywords:
3D Gaussians View Synthesis Real-Time Rendering
@@ 1 Introduction
Figure 1: Rendering results on the Waymo dataset [51]. Our method produces high-quality rendering at 135 FPS (1066$\times$1600) within half an hour of training. Current SOTA methods [64, 69] suffer from high training and rendering cost.
@@
With the development of neural scene representations, there have been some methods [32, 52, 43, 38, 78] that attempt to reconstruct street scenes with neural radiance fields [35]. To improve the modeling capability, Block-NeRF [52] divides the scene into several blocks and represents each one with a NeRF network. Although this strategy enables photo-realistic rendering of large-scale street scenes, Block-NeRF suffers from long training time due to the large amount of network parameters. Moreover, it cannot handle dynamic vehicles on the street, which are crucial aspects in autonomous driving environment simulation.
@@
Recently, some methods [39, 71, 64, 23] propose to represent dynamic driving scenes as compositional neural representations that consist of foreground moving cars and static background. To handle the dynamic car, they leverage tracked vehicle poses to establish the mapping between the observation space and the canonical space, where they use NeRF networks to model the car’s geometry and appearance. Although these methods produce reasonable results, they are still limited to the high training cost and low rendering speed.
@@
In this work, we propose a novel explicit scene representation for reconstructing dynamic 3D street scenes from images. The basic idea is utilizing point clouds to build dynamic scenes, which significantly increases the training and rendering efficiency. Specifically, we decompose urban street scenes into the static background and moving vehicles, which are separately built based on 3D Gaussians [19]. To handle the dynamics of foreground vehicles, we model their geometry as a set of points with optimizable tracked vehicle poses, where each point stores learnable 3D Gaussian parameters. Furthermore, the time-varying appearance is represented by a 4D spherical harmonics model that uses a time series function to predict spherical harmonics coefficients at any time step. Thanks to the dynamic 3D Gaussians representation, we can faithfully reconstruct the target urban street within half an hour and achieve real-time rendering (135FPS@1066x1600). Building upon the proposed scene representation, we develop several strategies to further improve the rendering performance, including the tracked pose optimization,point cloud initialization, and sky modeling.
@@
We evaluate the proposed method on Waymo Open [51] (Waymo) and KITTI [15] datasets, which present dynamic street scenes with complex vehicle motions and various environment conditions. Across all datasets, our approach achieves state-of-the-art performance in terms of rendering quality, while being rendered over 100 times faster than previous methods [64, 39, 69]. Furthermore, detailed ablations and scene editing applications are conducted to demonstrate the effectiveness of proposed components and the flexibility of the proposed representation, respectively.
@@
  • We propose several strategies including 4D spherical harmonics appearance model, tracked pose optimization, and point cloud initialization, which largely improve the rendering performance of Street Gaussians.
@@
2 Related work
Static scene modeling.  Neural scene representation proposes to represent 3D scenes with neural networks, which can model complex scenes from images through differentiable rendering. NeRF [35, 3, 4, 5, 36] represents continuous volumetric scenes with MLP networks and achieves impressive rendering results. Some works have been proposed to extend NeRF to urban scenes [43, 52, 16, 32, 30, 38, 55, 18, 9]. GridNeRF [66] proposes muti-resolution feature planes to help NeRF generate photorealistic results on large-scale scenes. DNMP [32] models the scene with deformable mesh primitives initialized by voxelizing point clouds. NeuRas [30] takes scaffold mesh as input and optimizes the neural texture field to perform fast rasterization.
Point-based rendering works [45, 22, 1, 10, 27] define learned neural descriptors on point clouds and perform differentiable rasterization with a neural renderer. However, they require dense point clouds as input and generate blurry results under regions with low point counts. A very recent work 3D Gaussian Splatting (3D GS) [19] defines a set of anisotropic Gaussians in 3D world and performs adaptive density control to achieve high-quality rendering results with only sparse point clouds input. However, 3D GS assumes the scene to be static and can not model dynamic moving objects.
Dynamic scene modeling.  Recent methods build 4D neural scene representation on single-object scenes by encoding time as additional input [50, 40, 13, 26, 2, 29, 41, 28]. Some works learn a scene decomposition of outdoor scenes under the supervision of optical flow [56] or vision transformer feature [69]. However, their scene representation is not instance aware, limiting the applications for autonomous driving simulation. Another line of works model the scene as the composition of moving object models and a background model [39, 23, 64, 71, 65, 54] with neural fields, which is most similar to us. However, they suffer from high memory cost on large scale scene and can not perform real-time rendering.
Extending point-based rendering to dynamic scene is also investigated recently [68, 76]. Recent approaches extend 3D GS to small-scale dynamic scenes by introducing deformation field [75, 61], physical priors [33] or 4D parametrization [73] to 3D Gaussian model. More recently, some concurrent works [8, 79] also explore 3D Gaussians in urban street scenes. DrivingGaussian [79] introduces Incremental 3D Static Gaussians and Composite Dynamic Gaussian Graphs. PVG [8] utilizes Periodic Vibration 3D Gaussians to model dynamic urban scene.
Simulation environments for autonomous driving.  Existing self-driving simulation engines such as CARLA [11] or AirSim [47] suffer from costly manual effort to create virtual environments and the lack of realism in the generated data. In recent years, a lot of effort has been put into building sensor simulations from autonomous driving data captured in real scenes. Some works [34, 74, 12] concentrate on LiDAR simulation by aggregating LiDAR and reconstructing textured primitives. However, they have difficulty handling high-resolution images and usually produce noisy appearance. Other works [72, 7, 57] reconstruct objects from multi-view images and LiDAR input, which can be interacted with other environments. However, these methods are restricted to existing images and fail to render novel views. Some methods utilize neural fields to perform multiply tasks including view synthesis [71, 39, 17], perception [23, 14, 78], generation [48, 37, 67, 70, 25] and inverse rendering [58, 60, 59, 42] on driving scenes. However, they struggle with high training and rendering cost. In contrast, Our method focuses on performing real-time rendering of dynamic urban scenes, which is crucial for autonomous driving simulation.
3 Method
Given a sequence of images captured from a moving vehicle in an urban street scene, our goal is to develop a model capable of generating photorealistic images for view synthesis. Towards this objective, we propose a novel scene representation, named Street Gaussians, specifically designed for representing dynamic street scenes. As shown in the Figure 2, we represent a dynamic urban street scene as a set of point clouds, each corresponding to either the static background or a moving vehicle (Section 3.1). The explicit point-based representation allows easy composition of separate models, enabling real-time rendering as well as the decomposition of foreground objects for editing applications (Section 3.2). The proposed scene representation can be effectively trained along with tracked vehicle poses from an off-the-shelf tracker, enhanced by our pose optimization strategy (Section 3.3).
@@ 3.1 Street Gaussians
Background model.  The background model is represented as a set of points in the world coordinate system. Each point is assigned with a 3D Gaussian to softly represent the continuous scene geometry and color. The Gaussian parameters consist of a covariance matrix $\bm{\Sigma}_{b}$ and a position vector $\bm{\mu}_{b}\in\mathbb{R}^{3}$, which denotes the mean value. To avoid invalid value during optimization, each covariance matrix is further reduced to a scaling matrix $\mathbf{S}_{b}$ and a rotation matrix $\mathbf{R}_{b}$, where $\mathbf{S}_{b}$ is characterized by its diagonal elements and $\mathbf{R}_{b}$ is converted into a unit quaternion.The covariance matrix $\bm{\Sigma}_{b}$ can be recovered from $\mathbf{S}_{b}$ and $\mathbf{R}_{b}$ as:
@@
Object model.  Consider a scene containing $N$ moving foreground object vehicles. Each object is represented with a set of optimizable tracked vehicle poses and a point cloud, where each point is assigned a 3D Gaussian, semantic logits, and a dynamic appearance model.
@@
$\mathbf{R}_{w}$ $=\mathbf{R}_{t}\mathbf{R}_{o},$
@@
The semantic representation of the object model is different from that of background. The main difference is that the semantic of the object model is a learnable one-dimensional scalar $\beta_{o}$ which represents the vehicle semantic class from the trakcer instead of a $M$-dimensional vector $\bm{\beta}_{b}$.
Initialization.  The SfM [46] point cloud used in 3D Gaussian is suitable for object centric scene. However, it can not provide good initialization for urban street scenes with many under-observed or textureless regions. We instead use aggregated LiDAR point cloud captured by ego vehicle as initialization. The colors of LiDAR point cloud are obtained by projecting to the corresponding image plane and querying the pixel value.
To initialize the object model, we first collect aggregated points inside the 3D bounding boxes and transform them into the local coordinate system. For object with less than 2K LiDAR points, we instead randomly sample 8K points inside the 3D bounding box as initialization. For the background model, we perform voxel downsampling for the remaining point cloud and filter out the points which are invisible to the training cameras. We incorporate SfM point cloud to compensate for the limited coverage of LiDAR over large areas.
@@ 3.2 Rendering of Street Gaussians
To render Street Gaussians, we need to aggregate the contribution of each model to render the final image. Previous methods [39, 23, 71, 64] require compositional rendering with complex raymarching because of neural field representation. Instead, Street Gaussians can be rendered by contacting all the point clouds and projecting them to 2D image space. Specifically, given a rendered time step $t$, we first compute spherical harmonics Eq. 3 , and transform the object point cloud into the world coordinate system using and Eq. 2 according to tracked vehicle pose $(\mathbf{R}_{t},\mathbf{T}_{t})$. Then we concatenate the background point cloud and the transformed object point clouds to form a new point cloud. To project this point cloud to 2D image space with camera extrinsic $\mathbf{W}$ and intrinsic $\mathbf{K}$, we compute the 2D Gaussian for each point in the point cloud [80]:
@@
where $\mathbf{J}$ is the Jacobian matrix of $\mathbf{K}$ $\mu^{\prime}$ and $\Sigma^{\prime}$ are the position and covariance matrix in 2D image space, respectively. Point-based $\alpha$-blending for each pixel is used to compute the color $\mathbf{C}$:
@@
Here $\alpha_{i}$ is the opacity $\alpha$ multiplied by the probability of the 2D Gaussian and $\mathbf{c}_{i}$ is the color computed from spherical harmonics $\mathbf{z}$ with the view direction. We can also render other signals like depth, opacity and semantic. For instance, the semantic map is rendered by changing color $c$ in Eq. 5 to semantic logits $\bm{\beta}$.
Since 3D Gaussian is defined in Euclidean space, it is inappropriate for them to model distant regions like sky. As a result, we utilize a high resolution cubemap which maps the view direction to sky color $\mathbf{C}_{\text{sky}}$. The explicit cubemap representation helps us recover details in sky regions without sacrificing inference speed. The final rendering color is obtained by blending $\mathbf{C}_{\text{sky}}$ and the color $\mathbf{C}$ in Eq. 5. More details can be found in the supplementary.
@@ 3.3 Training
Tracking Pose Optimization.  Positions and covariance matrices of the object Gaussians during rendering in Section 3.2 are closely correlated with the tracked pose parameters as shown in Eq. 2. However, bounding boxes produced by the tracker model are generally noisy. Directly using them to optimize our scene representation leads to degradation in rendering quality. As a result, we treat tracked poses as learnable parameters by adding a learnable transformation to each transformation matrix. Specifically, $\mathbf{R}_{t}$ and $\mathbf{T}_{t}$ in Eq. 2 are replaced by $\mathbf{R}_{t}^{{}^{\prime}}$ and $\mathbf{T}_{t}^{{}^{\prime}}$ which are defined as:
@@
Loss function.  We jointly optimize our scene representation, sky cubemap and tracked poses using the following loss function:
$\mathcal{L}=\mathcal{L}_{\text{color}}+\lambda_{1}\mathcal{L}_{\text{depth}}+\lambda_{2}\mathcal{L}_{\text{sky}}+\lambda_{3}\mathcal{L}_{\text{sem}}+\lambda_{4}\mathcal{L}_{\text{reg}}.$ (7)
In Eq. 9, $\mathcal{L}_{\text{color}}$ is the reconstruction loss between rendered and observed images following [19]. $\mathcal{L}_{\text{depth}}$ is a L1 loss between rendered depth and the depth generated by projecting sparse LiDAR points onto the camera plane. $\mathcal{L}_{\text{sky}}$ is a binary cross entropy loss for sky supervision. $\mathcal{L}_{\text{sem}}$ is an optional per-pixel softmax-cross-entropy loss between rendered semantic logits and input 2D semantic segmentation predictions [24] and $\mathcal{L}_{\text{reg}}$ is an regularization term used to remove floaters and enhance decomposition effects. Please refer to the supplementary material for details of each loss term.